A tank can be filled using pipes A, B, or both. It takes pipe A, running alone, 18 hours to fill the tank. It takes both pipes, running together, 9.9 hours to fill the tank. How long does it take pipe B, running alone, to fill the tank?
Answer by stanbon(75887)

It takes pipe A, running alone, 18 hours to fill the tank. It takes both pipes, running together, 9.9 hours to fill the tank. How long does it take pipe B, running alone, to fill the tank?
-----
A DATA:
time = 18 hrs/job ; rate = 1/18 jobs/hr
--------------------------------------------
A/B DATA:
time = 9.9 hrs/job ; rate = 1/9.9 job/hr
--------------------------------------------
B DATA:
time = x hrs/job ; rate = 1/x job/hr
----------------------------------------
Equation:
rate + rate = together rate
1/18 + 1/x = 1/9.9
---
9.9x + 9.9*18 = 18x
8.1x = 9.9*18
x = 22 hrs (time required by the B pipe to fill the tank)
===================
Cheers,
Stan H.